OpenStudy (ny,ny):

Solve each system by elimination 4x-9y=61 10x+3y=25

OpenStudy (espex):

If you multiply the bottom equation by 3, then add the two equation together, you will remove y and can solve for x, then plug x in and solve for y.

OpenStudy (ny,ny):

i get a really big decimal for x.

OpenStudy (espex):

Care to share your equations, something isn't right.

OpenStudy (ny,ny):

30x+9y=61 4x-9y=61 34x=86 x=?

OpenStudy (ny,ny):

ohhhh, i see my mistake. x=4?

OpenStudy (espex):

\[[10x + 3y = 25] \times 3 \rightarrow 30x + 9y = 75\] 4x - 9y = 61 +30x + 9y = 75

OpenStudy (espex):

Yep you got it.

OpenStudy (ny,ny):

thanks.

OpenStudy (espex):

You're welcome.